# RTNinja: a generalized machine learning framework for analyzing random telegraph noise signals in nanoelectronic devices

### v1

@@ Introduction
A comprehensive evaluation of the RTNinja framework’s performance across varying data complexities is presented in Figure 7 for three noise levels: 1%, 5%, and 30%. Additional results corresponding to intermediate noise levels of 10% and 20% are provided in Supplementary Figure 2. Figure 7 illustrates the comparison between the true number of sources in each dataset and the number estimated by RTNinja. The diagonal entries in this matrix indicate the number of datasets with correctly identified source counts. As noise levels increase, accurate detection becomes progressively more challenging due to weaker sources — i.e., those with low amplitudes — are increasingly suppressed by noise, often resulting in an underestimation of the true number of sources. Ideally, each row in the matrix should sum to 200 datasets, representing the complete batch for each source count. However, in practice, RTNinja occasionally fails to return a result for certain datasets (Supplementary Table 1). This is particularly true for datasets with fewer sources, where all sources become indistinguishable from noise due to low amplitudes. Conversely, in datasets with a higher number of sources and low noise levels, overfitting may occur due to the Bayesian algorithm overestimating the number of levels, which can lead to non-convergence in the SourcesMapper module. This is often caused by the initialization of the $N$-source hypothesis at an inflated value and the resulting computational burden of evaluating a large combinatorial space of candidate solutions. Hence, more datasets yield results at the highest noise level (30%) than at the lowest (1%) because the framework avoids overfitting in noisier conditions. Overall statistics on output yield across noise levels is provided in Supplementary Table 1.
@@
Figure 7: Performance of the RTNinja framework at 1%, 5%, and 30% noise levels across all dataset complexities. (a) Matrix showing the number of datasets corresponding to each combination of true and estimated source counts, with diagonal entries indicating correct matches. (b) Activity match accuracy for mapped input-output sources, grouped by dataset complexity. The number of detected sources and probability of achieving >95% activity match is indicated. (c) Accurate amplitude extraction is demonstrated, independent of noise level and data complexity. Blue squares indicate noise bounds; dashed lines mark 0.5×0.5$\times$0.5 × and 2×2$\times$2 × amplitude thresholds, with percentages showing sources within bounds. (d) Off-state and (e)On-state mean durations, with percentage of estimates within 0.5×0.5$\times$0.5 ×--2×2$\times$2 × of true values. Accuracy is affected by inherent statistical sampling limitations and increasing noise, which can lead to underestimation.
@@
Once the input-output source pairs have been successfully mapped, we assess RTNinja’s ability to accurately extract key characteristics of each individual source. The primary characteristic evaluated is the full activity profile of a source, measured using the activity match accuracy. To quantify this, we employ the Hamming distance, which calculates the dissimilarity between the binary activity patterns of the input and output, specifically whether a source is in the off- or on-state at each instance [34]. The Hamming distance is then converted into an activity match percentage by computing its complement. Figure 7 illustrates the distribution of activity match percentages plotted against cumulative probability for each $N$-source dataset configuration across varying noise levels. Overall, RTNinja exhibits strong performance in recovering source activity. However, the accuracy of activity matching systematically decreases with increasing data complexity, and also declines with higher noise levels for any fixed number of sources. These trends highlight the increasing challenge of accurately capturing source-level dynamics under more complex and noisier signal conditions.
@@
Figure 7 presents the comparison of input and output amplitudes for mapped sources across different $N$-source dataset types. To evaluate the accuracy of amplitude extraction, we compute the percentage of mapped sources whose estimated amplitudes fall within a factor of two (i.e., between 0.5×0.5$\times$0.5 × and 2×2$\times$2 ×) of the corresponding true amplitude. RTNinja demonstrates strong performance in accurately recovering source amplitudes, independent of noise levels. However, the overall detection window narrows as noise increases, resulting in a systematic reduction in the proportion of correctly identified sources — from 73.8% at 1% noise to 44.9% at 30% noise, as detailed in Supplementary Table 1. This drop is primarily attributed to low-amplitude sources being increasingly suppressed by the rising noise floor. Notably, the accuracy of amplitude estimation does not exhibit a clear dependency on data complexity, suggesting that noise, rather than the number of sources, is the dominant factor affecting amplitude detectability.
@@
Since the full binary activity trace of each source is available, we further analyze the durations of isolated on-state and off-state intervals. The temporal lengths of these continuous state segments approximate an exponential distribution, as discussed in Figure 2(a). We fit these distributions using a maximum likelihood estimation method to extract the primary parameters: the mean on-state duration ($\langle\tau_{off}^{n}\rangle$) and off-state duration ($\langle\tau_{on}^{n}\rangle$). The comparisons of these parameters between the ground truth and RTNinja outputs are shown in Figures 7 and 7, respectively. RTNinja successfully recovers these distribution parameters for a substantial proportion of sources, as assessed by the percentage of mapped sources for which the extracted mean lies within a factor of two of the true value. However, accuracy declines systematically with increasing noise, accompanied by a growing tendency to underestimate the mean durations. This underestimation arises when some sources are not detected, causing their activity to be erroneously merged into the trajectories of detected sources — resulting in artificially faster dynamics. Additionally, the reliability of mean duration estimates is inherently limited by the number of observed transitions: for example, if a source enters the on-state only twice during the observation window, the estimated mean is constrained by statistical sampling limitations.
@@
Figure 8: Overall impact of noise level on RTNinja performance. (a) Histogram showing the frequency distribution of the difference between the estimated and actual number of sources across all datasets, grouped by noise level. A progressive leftward shift illustrates increasing underestimation as noise rises. (b) Cumulative probability of source activity match accuracy, based on amplitude-centric input-output mapping. The total number of mapped sources and the fraction with >95% activity match is provided. (c) Complementary cumulative distribution function (CCDF) of estimated source amplitudes. Input (true) amplitude distributions are shown in grey as a reference. (d–e) Cumulative distributions of off-state and (e) on-state mean durations. True distributions are shown in grey. Deviations at higher noise levels reflect both missed sources and statistical compression due to limited sampling windows and temporal resolution.
@@
Figure 8(a) presents the cumulative frequency of the difference between the estimated and actual number of sources across noise levels. As anticipated, a systematic shift toward underestimation is observed with increasing noise, reflecting a higher fraction of missed sources as the noise floor rises. This trend is further supported by the activity match results shown in Figure 8(b), where activity accuracy — measured using the amplitude-based source matching strategy — declines inversely with noise. The primary source of this degradation lies in undetected sources whose activities are subsumed into the trajectories of detected ones. The overall distributions of mapped source characteristics — namely amplitudes, off-state mean durations, and on-state mean durations — are summarized in Figures 8(c)8(e). RTNinja consistently recovers the underlying amplitude distributions with high accuracy, regardless of noise level. However, the total number of detected sources systematically declines as noise increases, particularly due to the suppression of low-amplitude components that fall below the elevated noise floor. At the other extreme, a minor deviation in amplitude estimation is occasionally observed for very high-amplitude sources. For both off-state ($\langle\tau_{off}^{n}\rangle$) and on-state ($\langle\tau_{on}^{n}\rangle$) mean durations, a clear trend of underestimation emerges with rising noise levels. This is primarily caused by the merging of undetected source activity into the traces of detected sources, leading to artificially faster dynamics. Additionally, both the lower and upper tails of the extracted duration distributions are truncated across all noise conditions — a limitation imposed by the finite resolution and observation window of the RTN signal.
Conclusion
@@
Supplementary Information
Noise Level: 10% 20%
(a) Noise Level: 1%
(b) 5%
(c) 10%
(d) 20%
(e) 30%
Supplementary Figure 3: Difference between the number of sources estimated by RTNinja and the ground truth, shown across noise levels and grouped by $N$-source dataset type. A systematic shift toward negative differences is observed, indicating an increasing fraction of undetected sources as the noise level rises.
Noise Level (%) #Datasets with RTNinja Convergence #Datasets Yield (%) #Detected Sources #Detected Sources Yield(%)

### v2

@@ IV Performance Evaluation
A comprehensive evaluation of the RTNinja framework’s performance across varying data complexities is presented in Figure 4(e) for three noise levels: 1%, 5%, and 30%. Additional results corresponding to intermediate noise levels of 10% and 20% are provided in Supplementary Figure 2. Figure LABEL:fig_confmatrix illustrates the comparison between the true number of sources in each dataset and the number estimated by RTNinja. The diagonal entries in this matrix indicate the number of datasets with correctly identified source counts. As noise levels increase, accurate detection becomes progressively more challenging due to weaker sources — i.e., those with low amplitudes — are increasingly suppressed by noise, often resulting in an underestimation of the true number of sources. Ideally, each row in the matrix should sum to 200 datasets, representing the complete batch for each source count. However, in practice, RTNinja occasionally fails to return a result for certain datasets (Supplementary Table 1). This is particularly true for datasets with fewer sources, where all sources become indistinguishable from noise due to low amplitudes. Conversely, in datasets with a higher number of sources and low noise levels, overfitting may occur due to the Bayesian algorithm overestimating the number of levels, which can lead to non-convergence in the SourcesMapper module. This is often caused by the initialization of the $N$-source hypothesis at an inflated value and the resulting computational burden of evaluating a large combinatorial space of candidate solutions. Hence, more datasets yield results at the highest noise level (30%) than at the lowest (1%) because the framework avoids overfitting in noisier conditions. Overall statistics on output yield across noise levels is provided in Supplementary Table 1.
@@
(e) Performance of the RTNinja framework at 1%, 5%, and 30% noise levels across all dataset complexities. (a) Matrix showing the number of datasets corresponding to each combination of true and estimated source counts, with diagonal entries indicating correct matches. (b) Activity match accuracy for mapped input-output sources, grouped by dataset complexity. The number of detected sources and probability of achieving >>95% activity match is indicated. (c) Accurate amplitude extraction is demonstrated, independent of noise level and data complexity. Blue squares indicate noise bounds; dashed lines mark $0.5\times$ and $2\times$ amplitude thresholds, with percentages showing sources within bounds. (d) Off-state and (e)On-state mean durations, with percentage of estimates within $0.5\times$-$2\times$ of true values. Accuracy is affected by inherent statistical sampling limitations and increasing noise, which can lead to underestimation.
@@
Once the input-output source pairs have been successfully mapped, we assess RTNinja’s ability to accurately extract key characteristics of each individual source. The primary characteristic evaluated is the full activity profile of a source, measured using the activity match accuracy. To quantify this, we employ the Hamming distance, which calculates the dissimilarity between the binary activity patterns of the input and output, specifically whether a source is in the off- or on-state at each instance [34]. The Hamming distance is then converted into an activity match percentage by computing its complement. Figure LABEL:fig_gof illustrates the distribution of activity match percentages plotted against cumulative probability for each $N$-source dataset configuration across varying noise levels. Overall, RTNinja exhibits strong performance in recovering source activity. However, the accuracy of activity matching systematically decreases with increasing data complexity, and also declines with higher noise levels for any fixed number of sources. These trends highlight the increasing challenge of accurately capturing source-level dynamics under more complex and noisier signal conditions.
@@
Figure LABEL:fig_ampd presents the comparison of input and output amplitudes for mapped sources across different $N$-source dataset types. To evaluate the accuracy of amplitude extraction, we compute the percentage of mapped sources whose estimated amplitudes fall within a factor of two (i.e., between $0.5\times$ and $2\times$) of the corresponding true amplitude. RTNinja demonstrates strong performance in accurately recovering source amplitudes, independent of noise levels. However, the overall detection window narrows as noise increases, resulting in a systematic reduction in the proportion of correctly identified sources — from 73.8% at 1% noise to 44.9% at 30% noise, as detailed in Supplementary Table 1. This drop is primarily attributed to low-amplitude sources being increasingly suppressed by the rising noise floor. Notably, the accuracy of amplitude estimation does not exhibit a clear dependency on data complexity, suggesting that noise, rather than the number of sources, is the dominant factor affecting amplitude detectability.
@@
Since the full binary activity trace of each source is available, we further analyze the durations of isolated on-state and off-state intervals. The temporal lengths of these continuous state segments approximate an exponential distribution, as discussed in Figure LABEL:fig_mcsim. We fit these distributions using a maximum likelihood estimation method to extract the primary parameters: the mean on-state duration ($\langle\tau_{off}^{n}\rangle$) and off-state duration ($\langle\tau_{on}^{n}\rangle$). The comparisons of these parameters between the ground truth and RTNinja outputs are shown in Figures LABEL:fig_toff and LABEL:fig_ton, respectively. RTNinja successfully recovers these distribution parameters for a substantial proportion of sources, as assessed by the percentage of mapped sources for which the extracted mean lies within a factor of two of the true value. However, accuracy declines systematically with increasing noise, accompanied by a growing tendency to underestimate the mean durations. This underestimation arises when some sources are not detected, causing their activity to be erroneously merged into the trajectories of detected sources — resulting in artificially faster dynamics. Additionally, the reliability of mean duration estimates is inherently limited by the number of observed transitions: for example, if a source enters the on-state only twice during the observation window, the estimated mean is constrained by statistical sampling limitations.
@@
(f) Overall impact of noise level on RTNinja performance. (a) Histogram showing the frequency distribution of the difference between the estimated and actual number of sources across all datasets, grouped by noise level. A progressive leftward shift illustrates increasing underestimation as noise rises. (b) Cumulative probability of source activity match accuracy, based on amplitude-centric input-output mapping. The total number of mapped sources and the fraction with >>95% activity match is provided. (c) Complementary cumulative distribution function (CCDF) of estimated source amplitudes. Input (true) amplitude distributions are shown in grey as a reference. (d–e) Cumulative distributions of off-state and (e) on-state mean durations. True distributions are shown in grey. Deviations at higher noise levels reflect both missed sources and statistical compression due to limited sampling windows and temporal resolution.
@@
Figure LABEL:fig_ndiff presents the cumulative frequency of the difference between the estimated and actual number of sources across noise levels. As anticipated, a systematic shift toward underestimation is observed with increasing noise, reflecting a higher fraction of missed sources as the noise floor rises. This trend is further supported by the activity match results shown in Figure LABEL:fig_ngof, where activity accuracy — measured using the amplitude-based source matching strategy — declines inversely with noise. The primary source of this degradation lies in undetected sources whose activities are subsumed into the trajectories of detected ones. The overall distributions of mapped source characteristics — namely amplitudes, off-state mean durations, and on-state mean durations — are summarized in Figures LABEL:fig_nampdLABEL:fig_nton. RTNinja consistently recovers the underlying amplitude distributions with high accuracy, regardless of noise level. However, the total number of detected sources systematically declines as noise increases, particularly due to the suppression of low-amplitude components that fall below the elevated noise floor. At the other extreme, a minor deviation in amplitude estimation is occasionally observed for very high-amplitude sources. For both off-state ($\langle\tau_{off}^{n}\rangle$) and on-state ($\langle\tau_{on}^{n}\rangle$) mean durations, a clear trend of underestimation emerges with rising noise levels. This is primarily caused by the merging of undetected source activity into the traces of detected sources, leading to artificially faster dynamics. Additionally, both the lower and upper tails of the extracted duration distributions are truncated across all noise conditions — a limitation imposed by the finite resolution and observation window of the RTN signal.
V Scope of applicability
This section outlines the important conditions and constraints under which the RTNinja framework can be effectively utilized. While the framework is designed to be flexible and robust across a range of tasks, its performance and reliability are contingent upon specific assumptions and data characteristics. Understanding these constraints is critical for practitioners to determine whether the framework is appropriate for a given problem, and to anticipate potential limitations or challenges. The main guidelines that define the scope of applicability are as follows:
(i) Stationary signal: The framework assumes that the RTN signal remains stationary over the observation period. If device conditions drift over time or external interferences are present, the stationarity assumption may be violated, potentially compromising the accuracy of the analysis. To ensure reliable extraction of RTN parameters, such non-stationary components should be removed prior to applying the framework. A dedicated algorithm for drift correction has been recently reported and can be used in conjunction with this framework to enhance robustness [35].
(ii) Two-state sources: The framework is well-suited for scenarios where each source behaves as a two-state system, allowing accurate identification and mapping of source dynamics. In cases where sources exhibit multi-state behavior, the SourcesMapper module may interpret them as multiple independent two-state sources. Nevertheless, the LevelsExtractor module remains applicable, as it can still reliably identify discrete levels. This consideration is particularly relevant for studies involving individual defects.
(iii) Independent source activity: The framework assumes that each source operates independently, with no electrical or physical coupling between them. This assumption enables reliable extraction and mapping of individual source dynamics, as the model treats each signal component as arising from a distinct, non-interacting source. Consequently, the framework does not account for coupling effects or correlated switching behavior. In experimental scenarios where source interactions are suspected, the interpretations of extracted parameters should therefore be made with caution.
The RTNinja framework has already demonstrated its versatility and reliability through successful application to a variety of experimental RTN datasets across different device technologies. It has been effectively utilized for the analysis of nanoscale transistors and emerging device architectures, providing consistent and physically meaningful extraction of RTN parameters under diverse measurement conditions [36, 37, 38, 39]. These successful applications highlight the framework’s robustness and adaptability, while also validating the practical relevance of the assumptions outlined above. Collectively, they illustrate that within its defined scope, RTNinja serves as a powerful and generalizable tool for RTN signal analysis in advanced device technologies.
VI Conclusion
@@
Supplementary Material
Please refer to the supplementary material for additional details on the RTNinja framework. Supplementary Table 1 provides statistics on output yield across different noise levels. Supplementary Figure 1 gives additional information on the construction and clustering of the $P_{T}-\Delta$ space. Supplementary Figure 2 provides additional information on the framework’s performance at 10% and 20% noise levels. Supplementary Figure 3 gives additional information on the difference between estimated and true source counts across noise levels and dataset types.
Acknowledgements.
This work was part of the imec Industrial Affiliation Program. The authors thank Alexander Grill for helpful discussions during the RTNinja framework development.
Author Declarations
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
A.V. and R.D. conceptualized the work with assistance from C.M. A.V. generated the RTN data with support from R.D. A.V., R.D., and P.R. developed the RTNinja framework. A.V. performed the data analysis, prepared the figures, and wrote the manuscript, with support from C.M. and R.D. All authors contributed to the discussions and interpretation of the results.
Code and Data Availability
The synthetic datasets, RTNinja output files, the code for generating datasets, and the code for generating statistics from output files are available at 10.5281/zenodo.16750728. The code for RTNinja framework is currently part of a pending patent application (US 18/936,911; EP 23207819.6; CN 202411558787.3) and is therefore subject to legal restrictions.
Supplementary Material
Noise Level (%)
#Datasets with
RTNinja Convergence
#Datasets Yield (%) #Detected Sources
#Detected Sources
Yield(%)
@@ Acknowledgements.
Noise Level: 10% 20%
Supplementary Figure 3: Difference between the number of sources estimated by RTNinja and the ground truth, shown across noise levels and grouped by NN-source dataset type. A systematic shift toward negative differences is observed, indicating an increasing fraction of undetected sources as the noise level rises.